OpenStudy (anonymous):

HOW DO YOU DO THIS? Find an equivalent exponential expression without denominators. Keep the same base. 1/3^3

OpenStudy (id21):

\[(1/3)^{3}=3^{-3}\]

OpenStudy (radar):

\[3^{-3}= 1/3^{3}\]

OpenStudy (radar):

When a number is to a 0 power it is 1\[3^{0}=1\]Have you been taught that?

OpenStudy (radar):

Hopefully so. Then the rules of dividing a number with exponents is: \[x ^{a}\over x ^{b}\]=\[x ^{a-b}\] With the above in mind, look at the problem:\[(1/3)^{3}=(3^{0}/3^{1})^{3}=(3^{-1})^{3}=3^{-3}\]

OpenStudy (radar):

Your base in this case is 3.